# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v2

@@ C.2. A reduction from $\widehat{Q}$ to $\overline{Q}$
Definition C.7 (Interval version of a database instance $\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$).
@@
$\overline{R}_{\sigma_{i}}\stackrel{{\scriptstyle\text{def}}}{{=}}\{(\text{\sf seg}_{N}(z_{1}\circ\cdots\circ z_{i}),x_{1},\ldots,x_{m_{i}})\mid\\ (z_{1},\ldots,z_{i},x_{1},\ldots,x_{m_{i}})\in\widehat{R}_{\sigma_{i}}\}$
@@
Theorem C.9 (Implicit in (Khamis et al., 2022)).
@@
[…]
@@
Appendix D Missing Details from Section 7
Theorem 3.4 gives an upper bound on the amortized update time for both the $\text{IVM}^{\pm}$ and $\text{IVM}^{\pm}_{\delta}$ problems in terms of $N$, the length of the update stream. Corollary 3.5 tightens this upper bound from $N$ to the current size of the database, $|\mathcal{D}^{(\tau)}|$. In contrast, Theorem 3.6 gives a lower bound on the amortized update time only for $\text{IVM}^{\pm}_{\delta}$ (in terms of $N$ and by extension $|\mathcal{D}^{(\tau)}|$). We prove Theorem 3.4 for $\text{IVM}^{\pm}$ in Section D.1 and for $\text{IVM}^{\pm}_{\delta}$ in Section D.2. We prove Corollary 3.5 in Section D.3. In Section D.4, we prove Theorem 3.6 (for $\text{IVM}^{\pm}_{\delta}$).
D.1. Upper Bound for $\text{IVM}^{\pm}[Q]$
@@
Remark D.1 ().
If the number of updates $N$ is not known in advance, we can initialize $N$ to $1$ and keep doubling $N$ every time we reach $N$ updates. We show in the proof of Lemma D.3 that amortized analysis maintains the same overall runtime.
@@
  • $N$: The length of the update stream $\triangleright$ See Remark D.1
@@
  • $\widehat{\mathcal{D}}=\text{\sf CP}_{N}(\overline{\mathcal{D}})$$\triangleright$ Definition C.4
Insertion $\delta_{\tau}\mathcal{D}=\{+R_{j}(\bm{t})\}$normal-$\triangleright$ The $\tau$-th update inserts a tuple $\bm{t}$ into $R_{j}$ $\delta_{\tau}\overline{\mathcal{D}}\leftarrow\{+\overline{R}_{j}([\tau,\infty],\bm{t})\}$ normal-$\triangleright$ Insert $([\tau,\infty],\bm{t})$ into $\overline{R}_{j}$ $\overline{\mathcal{D}}\leftarrow\overline{\mathcal{D}}\biguplus\delta_{\tau}\overline{D}$normal-$\triangleright$ Apply $\delta_{\tau}\overline{\mathcal{D}}$ to $\overline{\mathcal{D}}$ $\delta_{\tau}\widehat{\mathcal{D}}\leftarrow\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ normal-$\triangleright$ Definition C.4 $\widehat{\mathcal{D}}\leftarrow\widehat{\mathcal{D}}\biguplus\delta_{\tau}\widehat{\mathcal{D}}$ normal-$\triangleright$ Use Theorem 3.1
Deletion $\delta_{\tau}\mathcal{D}=\{-R_{j}(\bm{t})\}$normal-$\triangleright$ The $\tau$-th update deletes a tuple $\bm{t}$ from $R_{j}$ Find $([\tau^{\prime},\infty],\bm{t})\in\overline{R}_{j}$ $\delta_{\tau}\overline{\mathcal{D}}\leftarrow\{-\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}),+\overline{R}_{j}([\tau^{\prime},\tau],\bm{t})\}$ normal-$\triangleright$ Replace $([\tau^{\prime},\infty],\bm{t})$ with $([\tau^{\prime},\tau],\bm{t})$ in $\overline{R}_{j}$ $\overline{\mathcal{D}}\leftarrow\overline{\mathcal{D}}\biguplus\delta_{\tau}\overline{D}$ $\delta_{\tau}\widehat{\mathcal{D}}\leftarrow\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ $\widehat{\mathcal{D}}\leftarrow\widehat{\mathcal{D}}\biguplus\delta_{\tau}\widehat{\mathcal{D}}$ normal-$\triangleright$ Proposition D.2
@@
For the rest of this section, we prove that Algorithm 1 meets Theorem 3.4. Fix a query $Q$ together with its time extension $\overline{Q}$ (Definition 6.2) and its multivariate extension $\widehat{Q}$ (Eq. (3)). We split the proof of Theorem 3.4 for $\text{IVM}^{\pm}[Q]$ into three parts. First, we reduce the maintenance of $Q$ to the maintenance of $\overline{Q}$ (Lemma D.1). Then, we reduce the latter problem to the maintenance of $\widehat{Q}$ (Lemma D.3). Finally, we show that $\widehat{Q}$ can be maintained with $\widetilde{\mathcal{O}}(N^{{\mathsf{w}(\widehat{Q})}-1})$ amortized update time, where ${\mathsf{w}(\widehat{Q})}$ is the fractional hypertree width of $\widehat{Q}$ (Lemma D.4). Theorem 3.4 for $\text{IVM}^{\pm}[Q]$ then follows from Lemmas D.1, D.3, and D.4.
D.1.1. From the Maintenance of $Q$ to the Maintenance of $\overline{Q}$
@@
Problem: $\overline{\text{IVM}}[\overline{Q}]$ Parameter: A time extension query $\overline{Q}$ Given: Restricted update stream $\delta_{1}\overline{\mathcal{D}},\delta_{2}\overline{\mathcal{D}},\ldots$ Task: Ensure constant-delay enumeration of $\overline{Q}_{\tau}(\overline{\mathcal{D}}^{(\tau)})$ after each update $\delta_{\tau}\overline{\mathcal{D}}$ We make the reduction from $\text{IVM}^{\pm}[Q]$ to $\overline{\text{IVM}}[\overline{Q}]$ precise:
Lemma D.1 ().
[…]
@@
D.1.2. From the Maintenance of $\overline{Q}$ to the Maintenance of $\widehat{Q}$
Now, we give a reduction from the maintenance of $\overline{Q}$ to the maintenance of $\widehat{Q}$. Let $\widehat{\mathcal{D}}$ be the database instance for $\widehat{Q}$. In order to translate an update $\delta_{\tau}\overline{\mathcal{D}}$ for $\overline{Q}$ into an update $\delta_{\tau}\widehat{\mathcal{D}}$ for $\widehat{Q}$, we take the canonical partition of $\delta_{\tau}\overline{\mathcal{D}}$, as given by Definition C.4. In particular, suppose that $N$, the length of the update stream, is a power of 2 and is known in advance. (See Remark D.1 about removing this assumption.) The corresponding update to $\widehat{Q}$ is
@@ D.1.2. From the Maintenance of $\overline{Q}$ to the Maintenance of $\widehat{Q}$
Lemma D.3 ().
[…]
@@
The reduction expressed in Lemma D.3 works as follows. Consider an update stream $\delta_{1}\overline{\mathcal{D}},\delta_{2}\overline{\mathcal{D}}\ldots$ for $\overline{Q}$. We transform each update $\delta_{\tau}\overline{\mathcal{D}}$ into its canonical partition $\delta_{\tau}\widehat{\mathcal{D}}$ with respect to $N$, and apply $\delta_{\tau}\widehat{\mathcal{D}}$ on the database $\widehat{\mathcal{D}}$ for $\widehat{Q}$. As a result, the algorithm maintains the invariant that the database $\widehat{\mathcal{D}}$ is always the canonical partition of the database $\overline{\mathcal{D}}$ with respect to $N$. Formally,
@@
Moreover, by Definition C.3, the number of single-tuple updates in $\delta_{\tau}\widehat{\mathcal{D}}$ is at most polylogarithmic in $N$ (which is $\widetilde{\mathcal{O}}(1)$).
Theorem C.5 connects the full outputs of $\overline{Q}(\overline{\mathcal{D}}^{(\tau)})$ and $\widehat{Q}(\widehat{\mathcal{D}}^{(\tau)})$ using the following relation:
@@
Finally, if the length of the update stream $N$ is not known in advance, then we use the doubling trick described in Remark D.1. In particular, we initialize $N$ to 1. Before we process $\tau$-th update for $\tau=N+1$, we replace the segment tree $\mathcal{T}_{N}$ by $\mathcal{T}_{2N}$ and replace $N$ by $2N$. Replacing the segment tree corresponds to creating a copy of the old tree and adding a new root node whose two children are the old segment tree and the copy. After updating the segment tree, we can redo the previous inserts into the new tree. The total update time would be
@@
D.1.3. Update Time for $\widehat{Q}$
@@
Lemma D.4 ().
[…]
@@
Recall from Appendix A that a tree decomposition $(\mathcal{T},\chi)$ of a query $Q$ is called optimal if it minimizes the fractional hypertree width of $Q$.
Proposition D.6 ().
@@
Proof of Proposition D.6.
The proof is based on the equivalence of tree decompositions and variable elimination orders. See Appendix A for some background. In particular, the proof works as follows. Pick some optimal variable elimination order $\bm{\mu}$. We inductively prove the following:
Claim D.2 ().
@@
Proof of Lemma D.4.
The proof of Lemma D.4 relies on the proof of Theorem 3.1 that is given in Section B.3. Consider an optimal tree decomposition $(\mathcal{T},\chi)$ for $\widehat{Q}$, i.e. where $\mathsf{w}(\mathcal{T},\chi)=\mathsf{w}(\widehat{Q})$. Let $\delta_{1}\widehat{\mathcal{D}},\delta_{2}\widehat{\mathcal{D}},\ldots$ be a restricted canonical update stream for $\widehat{Q}$. Recall that each $\delta_{\tau}\widehat{\mathcal{D}}$ is the canonical partition of a corresponding update $\delta_{\tau}\overline{\mathcal{D}}$ to $\overline{\mathcal{D}}$. Consider an update $\delta_{\tau}\widehat{\mathcal{D}}=\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ in this stream. First, assume that $\delta_{\tau}\overline{\mathcal{D}}$ is an insert, i.e. $\delta_{\tau}\overline{\mathcal{D}}=\{+\overline{R}_{j}([\tau,\infty],\bm{t}_{j})\}$ for some $j\in[k]$ and tuple $\bm{t}_{j}$ with schema $\bm{X}_{j}$. In this case, the corresponding $\delta_{\tau}\widehat{\mathcal{D}}=\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ is a set of single-tuple inserts, and we can apply these inserts to the materialized bags of the tree decomposition $(\mathcal{T},\chi)$, just as explained in the proof of Theorem 3.1. Truncation is more tricky. In particular, suppose that $\delta_{\tau}\overline{\mathcal{D}}$ is a truncation, i.e. $\delta_{\tau}\overline{\mathcal{D}}=\{-\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}_{j}),+\overline{R}_{j}([\tau^{\prime},\tau],\bm{t}_{j})\}$ for some $\tau^{\prime}<\tau$, $j\in[m]$ and tuple $\bm{t}_{j}$ over schema $\bm{X}_{j}$. In this case, we update the materialized bags of $(\mathcal{T},\chi)$ using the updates given by $\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$. Some of these updates are inserts, namely $\text{\sf CP}_{N}(\{+\overline{R}_{j}([\tau^{\prime},\tau],\bm{t}_{j})\})$, and these are accounted for by the analysis from the proof of Theorem 3.1. In particular, they are covered by the $\widetilde{\mathcal{O}}(N^{{\mathsf{w}(\widehat{Q})}})$ total update time over all inserts combined. However, some of the updates in $\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ are deletes, namely $\text{\sf CP}_{N}(\{-\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}_{j})\})$, and in order to account for those, we will use amortized analysis. In particular, we will deposit extra credits with every insert so that we can use them later to pay for the deletes during truncation. We explain this argument in detail below.
We start with the following claim, which is very similar in spirit to Proposition D.2:
Claim D.4 ().
@@
The above claim follows from Claim D.4 along with the fact that every bag in the tree decomposition $(\mathcal{T},\chi)$ is $Z$-prefix closed (Proposition D.6). In particular, Claim D.4 basically says that the two intervals $[\tau,\infty]$ and $[\tau,\tau]$ are “indistinguishable” as far as joins are concerned. Proposition D.6 ensures that whenever we take a projection in the query plan corresponding to the tree decomposition $(\mathcal{T},\chi)$, we are basically replacing these two intervals with two new intervals that contain them. Hence, the new intervals remain indistinguishable.
Now, we can analyze the total update time as follows. Claim D.5 implies that the total time needed to apply provisional inserts is upper bounded by the total time needed to apply finalized inserts. Hence, the overall runtime is dominated by finalized inserts. Moreover, every time we apply a provisional insert, we can deposit one credit with every tuple that we insert into any bag of $(\mathcal{T},\chi)$, in order to pay for the potential delete of this tuple in the future. These credits cover the cost of all deletes of provisional tuples. (Recall that finalized tuples cannot be deleted.) Finally, the total time needed to apply all finalized inserts is bounded by $\widetilde{\mathcal{O}}(N^{\mathsf{w}(\widehat{Q})})$, by Theorem 3.1.
@@
D.2. Upper Bound for $\text{IVM}^{\pm}_{\delta}[Q]$

### v3

@@ D.2. A reduction from $\widehat{Q}$ to $\overline{Q}$
Definition D.7 (Interval version of a database instance $\text{\sf IV}(\widehat{\mathcal{D}}_{\bm{\sigma}})$).
@@
$\overline{R}_{\sigma_{i}}\stackrel{{\scriptstyle\text{def}}}{{=}}\{(\text{\sf seg}_{N}(z_{1}\circ\cdots\circ z_{i}),x_{1},\ldots,x_{m_{i}})\mid(z_{1},\ldots,z_{i},x_{1},\ldots,x_{m_{i}})\in\widehat{R}_{\sigma_{i}}\}$
@@
Theorem D.9 (Implicit in (Khamis et al., 2022)).
@@
(44) $H_{k}([z],x_{1},\ldots,x_{m})\stackrel{{\scriptstyle\text{def}}}{{=}}\{(z_{1},\ldots,z_{k},x_{1},\ldots,x_{m})\mid(z_{1}\circ\cdots\circ z_{k})\in\text{\sf CP}_{N}([z])\wedge|z_{1}|=\cdots=|z_{k}|\}$
@@
Appendix E Missing Details from Section 7
Theorem 3.4 gives an upper bound on the amortized update time for both the $\text{IVM}^{\pm}$ and $\text{IVM}^{\pm}_{\delta}$ problems in terms of $N$, the length of the update stream. Corollary 3.5 tightens this upper bound from $N$ to the current size of the database, $|\mathcal{D}^{(\tau)}|$. In contrast, Theorem 3.6 gives a lower bound on the amortized update time only for $\text{IVM}^{\pm}_{\delta}$ (in terms of $N$ and by extension $|\mathcal{D}^{(\tau)}|$). We prove Theorem 3.4 for $\text{IVM}^{\pm}$ in Section E.1 and for $\text{IVM}^{\pm}_{\delta}$ in Section E.2. We prove Corollary 3.5 in Section E.3. In Section E.4, we prove Theorem 3.6 (for $\text{IVM}^{\pm}_{\delta}$).
E.1. Upper Bound for $\text{IVM}^{\pm}[Q]$
@@
Remark E.1.
If the number of updates $N$ is not known in advance, we can initialize $N$ to $1$ and keep doubling $N$ every time we reach $N$ updates. We show in the proof of Lemma E.3 that amortized analysis maintains the same overall runtime.
@@
  • $N$: The length of the update stream $\triangleright$ See Remark E.1
@@
  • $\widehat{\mathcal{D}}=\text{\sf CP}_{N}(\overline{\mathcal{D}})$$\triangleright$ Definition D.4
Insertion $\delta_{\tau}\mathcal{D}=\{+R_{j}(\bm{t})\}$$\triangleright$ The $\tau$-th update inserts a tuple $\bm{t}$ into $R_{j}$ $\delta_{\tau}\overline{\mathcal{D}}\leftarrow\{+\overline{R}_{j}([\tau,\infty],\bm{t})\}$triangleright Insert $([\tau,\infty],\bm{t})$ into $\overline{R}_{j}$ $\overline{\mathcal{D}}\leftarrow\overline{\mathcal{D}}\biguplus\delta_{\tau}\overline{D}$$\triangleright$ Apply $\delta_{\tau}\overline{\mathcal{D}}$ to $\overline{\mathcal{D}}$ $\delta_{\tau}\widehat{\mathcal{D}}\leftarrow\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$triangleright Definition D.4 $\widehat{\mathcal{D}}\leftarrow\widehat{\mathcal{D}}\biguplus\delta_{\tau}\widehat{\mathcal{D}}$triangleright Use Theorem 3.1
Deletion $\delta_{\tau}\mathcal{D}=\{-R_{j}(\bm{t})\}$$\triangleright$ The $\tau$-th update deletes a tuple $\bm{t}$ from $R_{j}$ Find $([\tau^{\prime},\infty],\bm{t})\in\overline{R}_{j}$ $\delta_{\tau}\overline{\mathcal{D}}\leftarrow\{-\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}),+\overline{R}_{j}([\tau^{\prime},\tau],\bm{t})\}$triangleright Replace $([\tau^{\prime},\infty],\bm{t})$ with $([\tau^{\prime},\tau],\bm{t})$ in $\overline{R}_{j}$ $\overline{\mathcal{D}}\leftarrow\overline{\mathcal{D}}\biguplus\delta_{\tau}\overline{D}$ $\delta_{\tau}\widehat{\mathcal{D}}\leftarrow\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ $\widehat{\mathcal{D}}\leftarrow\widehat{\mathcal{D}}\biguplus\delta_{\tau}\widehat{\mathcal{D}}$triangleright Proposition E.2
@@
For the rest of this section, we prove that Algorithm 2 meets Theorem 3.4. Fix a query $Q$ together with its time extension $\overline{Q}$ (Definition 6.2) and its multivariate extension $\widehat{Q}$ (Eq. (3)). We split the proof of Theorem 3.4 for $\text{IVM}^{\pm}[Q]$ into three parts. First, we reduce the maintenance of $Q$ to the maintenance of $\overline{Q}$ (Lemma E.1). Then, we reduce the latter problem to the maintenance of $\widehat{Q}$ (Lemma E.3). Finally, we show that $\widehat{Q}$ can be maintained with $\widetilde{\mathcal{O}}(N^{{\mathsf{w}(\widehat{Q})}-1})$ amortized update time, where ${\mathsf{w}(\widehat{Q})}$ is the fractional hypertree width of $\widehat{Q}$ (Lemma E.4). Theorem 3.4 for $\text{IVM}^{\pm}[Q]$ then follows from Lemmas E.1, E.3, and E.4.
E.1.1. From the Maintenance of $Q$ to the Maintenance of $\overline{Q}$
@@
Problem: $\overline{\text{IVM}}[\overline{Q}]$ Parameter: A time extension query $\overline{Q}$ Given: Restricted update stream $\delta_{1}\overline{\mathcal{D}},\delta_{2}\overline{\mathcal{D}},\ldots$ Task: Ensure constant-delay enumeration of $\overline{Q}_{\tau}(\overline{\mathcal{D}}^{(\tau)})$ after each update $\delta_{\tau}\overline{\mathcal{D}}$
We make the reduction from $\text{IVM}^{\pm}[Q]$ to $\overline{\text{IVM}}[\overline{Q}]$ precise:
Lemma E.1.
If $\overline{\text{IVM}}[\overline{Q}]$ can be solved with (amortized) $f(N)$ update time for any update stream of length $N$ and some function $f$, then $\text{IVM}^{\pm}[Q]$ can be solved with (amortized) $\mathcal{O}(f(N))$ update time.
@@
E.1.2. From the Maintenance of $\overline{Q}$ to the Maintenance of $\widehat{Q}$
Now, we give a reduction from the maintenance of $\overline{Q}$ to the maintenance of $\widehat{Q}$. Let $\widehat{\mathcal{D}}$ be the database instance for $\widehat{Q}$. In order to translate an update $\delta_{\tau}\overline{\mathcal{D}}$ for $\overline{Q}$ into an update $\delta_{\tau}\widehat{\mathcal{D}}$ for $\widehat{Q}$, we take the canonical partition of $\delta_{\tau}\overline{\mathcal{D}}$, as given by Definition D.4. In particular, suppose that $N$, the length of the update stream, is a power of 2 and is known in advance. (See Remark E.1 about removing this assumption.) The corresponding update to $\widehat{Q}$ is
@@ E.1.2. From the Maintenance of $\overline{Q}$ to the Maintenance of $\widehat{Q}$
Lemma E.3.
If $\widehat{\text{IVM}}[\widehat{Q}]$ can be solved with (amortized) $f(N)$ update time for any update stream of length $N$ and some monotonic function $f$, then $\overline{\text{IVM}}[\overline{Q}]$ can be solved with $\widetilde{\mathcal{O}}(f(N))$ amortized update time.
@@
The reduction expressed in Lemma E.3 works as follows. Consider an update stream $\delta_{1}\overline{\mathcal{D}},\delta_{2}\overline{\mathcal{D}}\ldots$ for $\overline{Q}$. We transform each update $\delta_{\tau}\overline{\mathcal{D}}$ into its canonical partition $\delta_{\tau}\widehat{\mathcal{D}}$ with respect to $N$, and apply $\delta_{\tau}\widehat{\mathcal{D}}$ on the database $\widehat{\mathcal{D}}$ for $\widehat{Q}$. As a result, the algorithm maintains the invariant that the database $\widehat{\mathcal{D}}$ is always the canonical partition of the database $\overline{\mathcal{D}}$ with respect to $N$. Formally,
@@
Moreover, by Definition D.3, the number of single-tuple updates in $\delta_{\tau}\widehat{\mathcal{D}}$ is at most polylogarithmic in $N$ (which is $\widetilde{\mathcal{O}}(1)$).
Theorem D.5 connects the full outputs of $\overline{Q}(\overline{\mathcal{D}}^{(\tau)})$ and $\widehat{Q}(\widehat{\mathcal{D}}^{(\tau)})$ using the following relation:
@@
Finally, if the length of the update stream $N$ is not known in advance, then we use the doubling trick described in Remark E.1. In particular, we initialize $N$ to 1. Before we process $\tau$-th update for $\tau=N+1$, we replace the segment tree $\mathcal{T}_{N}$ by $\mathcal{T}_{2N}$ and replace $N$ by $2N$. Replacing the segment tree corresponds to creating a copy of the old tree and adding a new root node whose two children are the old segment tree and the copy. After updating the segment tree, we can redo the previous inserts into the new tree. The total update time would be
@@
E.1.3. Update Time for $\widehat{Q}$
@@
Lemma E.4.
For any restricted canonical update stream of length $N$, the problem $\widehat{\text{IVM}}[\widehat{Q}]$ can be solved with $\widetilde{\mathcal{O}}(N^{{\mathsf{w}(\widehat{Q})}-1})$ amortized update time, where ${\mathsf{w}(\widehat{Q})}$ is the fractional hypertree width of $\widehat{Q}$.
@@
Recall from Appendix B that a tree decomposition $(\mathcal{T},\chi)$ of a query $Q$ is called optimal if it minimizes the fractional hypertree width of $Q$.
Proposition E.6.
@@
Proof of Proposition E.6.
The proof is based on the equivalence of tree decompositions and variable elimination orders. See Appendix B for some background. In particular, the proof works as follows. Pick some optimal variable elimination order $\bm{\mu}$. We inductively prove the following:
Claim E.2.
@@
Proof of Lemma E.4.
The proof of Lemma E.4 relies on the proof of Theorem 3.1 that is given in Section C.3. Consider an optimal tree decomposition $(\mathcal{T},\chi)$ for $\widehat{Q}$, i.e. where $\mathsf{w}(\mathcal{T},\chi)=\mathsf{w}(\widehat{Q})$. Let $\delta_{1}\widehat{\mathcal{D}},\delta_{2}\widehat{\mathcal{D}},\ldots$ be a restricted canonical update stream for $\widehat{Q}$. Recall that each $\delta_{\tau}\widehat{\mathcal{D}}$ is the canonical partition of a corresponding update $\delta_{\tau}\overline{\mathcal{D}}$ to $\overline{\mathcal{D}}$. Consider an update $\delta_{\tau}\widehat{\mathcal{D}}=\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ in this stream. First, assume that $\delta_{\tau}\overline{\mathcal{D}}$ is an insert, i.e. $\delta_{\tau}\overline{\mathcal{D}}=\{+\overline{R}_{j}([\tau,\infty],\bm{t}_{j})\}$ for some $j\in[k]$ and tuple $\bm{t}_{j}$ with schema $\bm{X}_{j}$. In this case, the corresponding $\delta_{\tau}\widehat{\mathcal{D}}=\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ is a set of single-tuple inserts, and we can apply these inserts to the materialized bags of the tree decomposition $(\mathcal{T},\chi)$, just as explained in the proof of Theorem 3.1. Truncation is more tricky. In particular, suppose that $\delta_{\tau}\overline{\mathcal{D}}$ is a truncation, i.e. $\delta_{\tau}\overline{\mathcal{D}}=\{-\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}_{j}),+\overline{R}_{j}([\tau^{\prime},\tau],\bm{t}_{j})\}$ for some $\tau^{\prime}<\tau$, $j\in[m]$ and tuple $\bm{t}_{j}$ over schema $\bm{X}_{j}$. In this case, we update the materialized bags of $(\mathcal{T},\chi)$ using the updates given by $\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$. Some of these updates are inserts, namely $\text{\sf CP}_{N}(\{+\overline{R}_{j}([\tau^{\prime},\tau],\bm{t}_{j})\})$, and these are accounted for by the analysis from the proof of Theorem 3.1. In particular, they are covered by the $\widetilde{\mathcal{O}}(N^{{\mathsf{w}(\widehat{Q})}})$ total update time over all inserts combined. However, some of the updates in $\text{\sf CP}_{N}(\delta_{\tau}\overline{\mathcal{D}})$ are deletes, namely $\text{\sf CP}_{N}(\{-\overline{R}_{j}([\tau^{\prime},\infty],\bm{t}_{j})\})$, and in order to account for those, we will use amortized analysis. In particular, we will deposit extra credits with every insert so that we can use them later to pay for the deletes during truncation. We explain this argument in detail below.
We start with the following claim, which is very similar in spirit to Proposition E.2:
Claim E.4.
@@
The above claim follows from Claim E.4 along with the fact that every bag in the tree decomposition $(\mathcal{T},\chi)$ is $Z$-prefix closed (Proposition E.6). In particular, Claim E.4 basically says that the two intervals $[\tau,\infty]$ and $[\tau,\tau]$ are “indistinguishable” as far as joins are concerned. Proposition E.6 ensures that whenever we take a projection in the query plan corresponding to the tree decomposition $(\mathcal{T},\chi)$, we are basically replacing these two intervals with two new intervals that contain them. Hence, the new intervals remain indistinguishable.
Now, we can analyze the total update time as follows. Claim E.5 implies that the total time needed to apply provisional inserts is upper bounded by the total time needed to apply finalized inserts. Hence, the overall runtime is dominated by finalized inserts. Moreover, every time we apply a provisional insert, we can deposit one credit with every tuple that we insert into any bag of $(\mathcal{T},\chi)$, in order to pay for the potential delete of this tuple in the future. These credits cover the cost of all deletes of provisional tuples. (Recall that finalized tuples cannot be deleted.) Finally, the total time needed to apply all finalized inserts is bounded by $\widetilde{\mathcal{O}}(N^{\mathsf{w}(\widehat{Q})})$, by Theorem 3.1.
@@
E.2. Upper Bound for $\text{IVM}^{\pm}_{\delta}[Q]$